### v1

@@ 3.1 Uniform Noisy Generatability
(of sufficiency in Theorem 3.1) Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose that $\Bigl{|}\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)\Bigl{|}=\infty.$ Then, the generator $\mathcal{G}$ which simply computes $S=\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)$ and plays arbitrarily from $S\setminus\{x_{1},\dots,x_{t}\}$ on every round $t\in\mathbb{N}$ generates perfectly (according to Definition A.1) from round $t=1$ itself, completing the proof. ∎
@@
(of necessity in Theorem 3.1) Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose that $\Bigl{|}\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)\Bigl{|}<\infty.$ We need to show that $\mathcal{H}$ is not uniformly noisily generatable. Let $\mathcal{F}\subseteq\mathcal{H}$ be the smallest subset of $\mathcal{H}$ such that
$\Bigl{|}\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)\Bigl{|}<\infty.$
Then, by definition, we know that for every $f\in\mathcal{F}$, we have that
$\Bigl{|}\bigcap_{h\in\mathcal{F}\setminus f}\operatorname{supp}(h)\Bigl{|}=\infty.$
We will show that $\mathcal{F}$ is not uniformly noisily generatable which will imply that $\mathcal{H}$ is not uniformly noisily generatable. To that end, we need to show that for every $\mathcal{G}$ and any $d\in\mathbb{N}$, there exists a $h\in\mathcal{F}$ and a sequence $x_{1},x_{2},\dots$ with
such that for every $t\in\mathbb{N}$ where |$|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h)|=d$, there exists an $s\geq t$ such that $\mathcal{G}(x_{1:s})\notin\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$. To that end, fix a generator $\mathcal{G}$ and a number $d\in\mathbbm{N}$. Fix some $f\in\mathcal{F}$ and let $x_{1},\dots,x_{d}$ be a sequence of $d$ unique examples in $\bigcap_{h\in\mathcal{F}\setminus f}\operatorname{supp}(h)$. Let $z_{1},\dots,z_{d}$ be any $d$ unique examples in $\operatorname{supp}(f)\setminus\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)$ and consider the sequence of $2d$ unique examples obtained by concatenation $z_{1},\dots,z_{d}$ to the end of $x_{1},\dots,x_{d}$. Denote this new sequence by $x_{1:2d}.$ Let $q=|\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)|$ and let $v_{1},\dots,v_{q}$ be its elements sorted in its natural order. Concatenate $v_{1},\dots,v_{q}$ to the end of $x_{1:2d}$ and denote this new sequence of unique examples as $x_{1:2d+q}.$ Observe that for every $h\in\mathcal{F}$, we have that $|\operatorname{supp}(h)\cap\{x_{1},\dots x_{2d+q}\}|\geq d.$ Let $\hat{x}_{2d+q}=\mathcal{G}(x_{1},\dots,x_{2d+q})$ and suppose without loss of generality that $\hat{x}_{2d+q}\notin\{x_{1},\dots,x_{2d+q}\}.$ Let $h^{\star}\in\mathcal{F}$ be a be hypothesis such that $\hat{x}_{2d+q}\notin\operatorname{supp}(h^{\star}).$ Such a hypothesis must exist because $\hat{x}_{2d+q}\notin\{x_{1},\dots,x_{2d+q}\}$ and $\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)\subseteq\{x_{1},\dots,x_{2d+q}\}.$ Let $x_{2d+q+1},x_{2d+q+2},\dots$ be any completion of the stream such that $\{x_{2d+q+t}\}_{t=1}^{\infty}\subseteq\operatorname{supp}(h^{\star})$ and $\{x_{2d+q+t}\}_{t=1}^{\infty}\cap\{x_{t}\}_{t=1}^{2d+q}=\emptyset.$
We are now ready to complete the proof. Let $h^{\star}$ and $x_{1},x_{2},\dots$ be the hypothesis and sequence chosen above. Then, by definition, observe that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h^{\star})\}=\sum_{t=1}^{2d+q}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h^{\star})\}\leq d<\infty.$
Moreover, we also know that $|\operatorname{supp}(h^{\star})\cap\{x_{1},\dots x_{2d+q}\}|\geq d$ and the first $2d+q$ examples are distinct, implying that there exists exactly one time point $t\leq 2d+q$ such that $|\operatorname{supp}(h^{\star})\cap\{x_{1},\dots x_{2d+q}\}|=d.$ Finally, noting that $2d+q\geq t$ and that $\mathcal{G}(x_{1},\dots,x_{2d+q})\notin\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{2d+1}\}$ completes the proof that $\mathcal{F}$ is not uniformly noisily generatable since $\mathcal{G}$ and $d$ were chosen arbitrarily. ∎
3.2 Noisy Uniform Generatability
Theorem 3.1 shows that obtaining guarantees that are uniform over the noise level is generally hopeless. In this section, we study the easier setting of noisy uniform generation, where the number of examples needed for perfect generation can depend on the noise level. At a high-level the results in this section extend the techniques from Li et al. (2024) to account for noisy example streams. As such, we first define a noisy analog of the Closure dimension.
Definition 3.1 ($n$-Noisy Closure dimension).
The Noisy Closure dimension of $\mathcal{H}$ at level $n\in\mathbbm{N}$, denoted $\operatorname{NC}_{n}(\mathcal{H})$, is the largest natural number $d$ for which there exists distinct $x_{1},\dots,x_{d}\in\mathcal{X}$ such that $\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},n}\neq\bot$ and $|\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},n}|<\infty.$ If this is true for arbitrarily large $d\in\mathbbm{N}$, then we say that $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ On the other hand, if this is not true for $d=1$, we say that $\operatorname{NC}_{n}(\mathcal{H})=0.$
Unlike the Closure dimension, the Noisy Closure dimension is scale-sensitive, it is defined with respect to every noise-level $n\in\mathbb{N}$. Scale-sensitive dimensions are not new to learning theory, but have also been defined and used to characterize other properties of hypothesis classes like PAC and online learnability for regression problems (Rakhlin et al., 2015; Bartlett et al., 1994). Our main theorem in this section uses the Noisy Closure dimension to provide a complete characterization of which classes are noisily uniform generatable.
Theorem 3.2 (Characterization of Noisy Uniform Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. The following statements are equivalent.
  • (i)
    $\mathcal{H}$ is noisily uniformly generatable.
  • (ii)
    $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for all $n\in\mathbb{N}.$
We start with the necessity direction.
Proof.
(of necessity in Theorem 3.2) Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose there exists a $n\in\mathbb{N}$ such that $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ We need to show that this means that $\mathcal{H}$ is not noisily uniformly generatable. In particular, we need to show that for every generator $\mathcal{G}$, there exists a noise level $n^{\star}\in\mathbb{N}$ such that for every $d\in\mathbbm{N}$, there exists a hypothesis $h\in\mathcal{H}$ and a sequence $x_{1},x_{2},\dots$ with
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star},$
so that for every $t\in\mathbbm{N}$ with |$|\{x_{1},\dots,x_{t}\}|=d$, there exists $s\geq t$ where $\mathcal{G}(x_{1:s})\notin\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$.To that end, let $\mathcal{G}$ be any generator and consider the noise level $n^{\star}=n$ so that $\operatorname{NC}_{n^{\star}}(\mathcal{H})=\infty.$ Fix any $d\in\mathbb{N}.$ We will now pick such a hypothesis $h^{\star}\in\mathcal{H}$ and a valid stream $x_{1},x_{2},\dots.$
Second, there exists only one $t\in\mathbb{N}$, namely $t=d$, such that |$|\{x_{1},\dots,x_{t}\}|=d$ as the first $d^{\star}+1>d$ examples of the stream are distinct by construction. However, observe that by construction, at time point $s=d^{\star}+q\geq d$, we have that $\mathcal{G}(x_{1},\dots,x_{s})=\hat{x}_{d^{\star}+q}\notin\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{s}\}$ by choice of $h^{\star}.$ Since $\mathcal{G}$ and $d\in\mathbb{N}$ were picked arbitrarily, this is true for all $\mathcal{G}$ and $d\in\mathbb{N}$, which completes the proof. ∎
Finally, we prove the sufficiency direction.
Proof.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. We need to show that if $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for all $n\in\mathbb{N},$ then $\mathcal{H}$ is noisily uniformly generatable. For a sequence $x_{1},x_{2},\dots$, let $d_{t}:=|\{x_{1},\dots,x_{t}\}|$ denote the number of distinct examples amongst the first $t$ examples. Consider the following generator $\mathcal{G}.$ At time $t\in\mathbb{N}$, $\mathcal{G}$ computes
$n_{t}:=\max\{n\in\mathbb{N}:\operatorname{NC}_{n}(\mathcal{H})<d_{t}\}\cup\{0\}.$
If $n_{t}=0,$ $\mathcal{G}$ plays any $x\in\mathcal{X}.$ Otherwise, it plays from $\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n_{t}}\setminus\{x_{1},x_{2},\dots,x_{t}\}.$ We now show that $\mathcal{G}$ is a noisy uniform generator. To that end, let $n^{\star}\in\mathbb{N}$ be the noise level chosen by the adversary. We claim that $\mathcal{G}$ can perfectly generate after observing $\operatorname{NC}_{n^{\star}}(\mathcal{H})+1$ distinct examples. To see why, let $h\in\mathcal{H}$ be any hypothesis, and $x_{1},x_{2},\dots$ be any stream such that
$\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq n^{\star}.$
Without loss of generality, suppose there exists $t^{\star}\in\mathbb{N}$ such that $d_{t^{\star}}=\operatorname{NC}_{n^{\star}}(\mathcal{H})+1.$ Fix any $s\geq t^{\star}$. By definition of $n_{t}$, it must be the case that $n_{s}\geq n^{\star}.$ Thus, $h\in\mathcal{H}(x_{1},\dots,x_{s},n_{s})$ and $\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H},n_{s}}\subseteq\operatorname{supp}(h)$. Moreover, because $d_{s}>\operatorname{NC}_{n_{s}}(\mathcal{H})$, again by definition, it must be the case that $|\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H},n_{s}}|=\infty.$ Accordingly, $\langle x_{1},\dots,x_{s}\rangle_{\mathcal{H},n_{s}}\setminus\{x_{1},\dots,x_{s}\}\neq\emptyset$ and $\mathcal{G}$ is guaranteed to output an example in $\operatorname{supp}(h)$ on round $s$. Since $s\geq t^{\star}$ was chosen arbitrarily, this is true for all such $s$, completing the proof. ∎
Unlike uniform noisy generation, our next result shows that all finite classes are noisily uniformly generatable.
Corollary 3.3 (All Finite Classes are Noisily Uniformly Generatable).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is finite, then $\mathcal{H}$ is noisily uniformly generatable.
Proof.
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be a finite hypothesis class with $q:=|\mathcal{H}|$ satisfying the $\operatorname{UUS}$ property. To show that $\mathcal{H}$ is noisily uniformly generatable, it suffices to show that $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for all $n\in\mathbb{N}.$ For any subset $V\subseteq\mathcal{H},$ define
$\langle\emptyset\rangle_{V}=\bigcap_{h\in V}\operatorname{supp}(h)\text{ and }\mathcal{F}=\{V\subseteq\mathcal{H}:|\langle\emptyset\rangle_{V}|<\infty\}$
to be the set of examples common to all hypotheses in $V$ and the subsets of $\mathcal{H}$ whose intersection of supports has finite cardinality, respectively. Then, let
$d:=\max_{V\in\mathcal{F}}|\langle\emptyset\rangle_{V}|,$
@@ 3.2 Noisy Uniform Generatability
be the maximum size of a finite set of examples common to a subset of $\mathcal{H}$. Note that $d$ is finite because $\mathcal{H}$ is finite. Fix any noise level $n\in\mathbb{N}$ and for the sake of contradiction, suppose $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ This means that for every $s\in\mathbb{N}$, there exists $t\geq s$ and a sequence of distinct examples ${x_{1},\dots,x_{t}}$ such that $|\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}|<\infty$. Pick $s=nq+d+1$ and consider the stream $x_{1},\dots,x_{t}$ such that $\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}<\infty$ for $t\geq s.$ Recall that $\mathcal{H}(x_{1:t},n)=\{h\in\mathcal{H}:|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h)|\geq t-n\}$ That is, $\mathcal{H}(x_{1:t},n)$ contains all hypotheses in $\mathcal{H}$ that are inconsistent with $x_{1:t}$ on at most $n$ examples. Since $|\mathcal{H}(x_{1:t},n)|\leq|\mathcal{H}|=q,$ there are at least
$t-nq\geq(nq+d+1)-nq=d+1$
distinct examples in $x_{1:t}$ contained in the support of all hypotheses in $\mathcal{H}(x_{1:t},n)$. In other words, $|\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}|=\langle\emptyset\rangle_{\mathcal{H}(x_{1:t},n)}\geq d+1>d.$ This is a contradiction as $d=\max_{V\in\mathcal{F}}|\langle\emptyset\rangle_{V}|$ by definition and $\mathcal{H}(x_{1:t},n)\in\mathcal{F}.$ Thus, $\operatorname{NC}_{n}(\mathcal{H})<nq+d+1<\infty$ for all $n\in\mathbb{N},$ completing the proof. ∎
@@
We end this section by establishing that noisy uniform generatability is strictly harder than noiseless uniform generatability. This is unlike PAC and online learnability for classification problems, where typically the noise-free (realizable) and noisy (agnostic) characterizations of learnability are equivalent (Littlestone, 1987; Ben-David et al., 2009).
Lemma 3.4 (Uniform Generatability $\neq$ Noisy Uniform Generatability).
@@
Let =$\mathbb{P}=\{p_{1},p_{2},\dots\}$ be the set of primes, $S_{q}=\{-q^{n}:n\in\mathbb{N}\}$, $\mathbb{S}=\bigcup_{q\in\mathbb{P}}S_{q}$ and $\mathcal{X}=\mathbb{P}\cup\mathbb{S}$. Let the support of each hypothesis $h_{p}$ be
$\operatorname{supp}(h_{p})=\mathbb{P}\setminus\{p\}\cup\mathbb{S}\setminus S_{p}$
and :=$\mathcal{H}:=\{h_{p}:p\in\mathbb{P}\}.$ Observe that $\mathcal{H}$ satisfies the $\operatorname{UUS}$ property and $\mathcal{X}$ is countable.
@@
We now prove that $\operatorname{C}(\mathcal{H})=0.$ Suppose the first element in the sequence is $x_{1}=-p^{n}$ for some $p\in\mathbb{P}$ and $n\in\mathbb{N}.$ Then $|\langle x_{1}\rangle_{\mathcal{H}}|=|\{-p^{m}:m\in\mathbb{N}\ \}|=\infty.$ If the first element in the sequence is $x_{1}=p\in\mathbb{P},$ then $|\langle x_{1}\rangle_{\mathcal{H}}|=|\{-p^{n}:n\in\mathbb{N}\ \}|=\infty.$ Note that these are the only two possible cases as it must be the case that $x_{1}\in\operatorname{supp}(h)$ for some $h\in\mathcal{H}.$
@@
Now, we show that $\operatorname{NC}_{1}(\mathcal{H})=\infty.$ Fix some $d\in\mathbb{N}.$ Consider the stream $p_{1},p_{2},\dots,p_{d}$ of the first $d$ prime numbers. Then, $|\langle p_{1},\dots,p_{d}\rangle_{\mathcal{H},1}|=|\emptyset|=0,$ due to the fact that $\mathcal{H}(p_{1},\dots,p_{d},1)=\mathcal{H}$ and the intersection of the supports of all $h\in\mathcal{H}$ is empty. So, $\operatorname{NC}_{1}(\mathcal{H})\geq d.$ Since $d\in\mathbb{N}$ was picked arbitrarily, this is true for all $d\in\mathbbm{N}$ and therefore $\mathcal{H}$ is not noisily uniformly generatable. ∎
@@
Lemma 3.4 shows a strong separation between noise-free and noisy uniform generation – there is a class that is uniformly generatable, but not noisily uniform generatable even when the adversary is only allowed to perturb at most one example.
3.3 Noisy Non-uniform Generatability

### v2

@@ 3.1 Uniform Noise-independent Generatability
(of necessity in Theorem 3.1) Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose that $\Bigl{|}\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)\Bigl{|}=:n<\infty.$ We need to show that for every $\mathcal{G}$ and sufficiently large $d\in\mathbb{N}$, there exists a $h\in\mathcal{H}$ and a sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty,$ such that for every $t\in\mathbb{N}$ where |$|\{x_{1},\dots,x_{t}\}|=d$, there exists an $s\geq t$ such that $\mathcal{G}(x_{1:s})\notin\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$. To that end, fix a generator $\mathcal{G}$ and a number $d\geq n$. Let $x_{1},\dots,x_{n}$ be the sequence of $n$ examples in $\bigcap_{h\in\mathcal{H}}\operatorname{supp}(h)$ sorted in their natural order. Pick any sequence of distinct examples $x_{n+1},x_{n+2},\dots,x_{d}$ and concatenate them to the end $x_{1:n}$. Let $\hat{x}=\mathcal{G}(x_{1:d})$ and suppose without loss of generality that $\hat{x}\notin\{x_{1:d}\}.$ Then, by construction, there exists a $h\in\mathcal{H}$ such that $\hat{x}\notin\operatorname{supp}(h).$ Finally, complete the stream by picking distinct $\{x_{d+1},x_{d+2},\dots\}\subseteq\operatorname{supp}(h).$ First, by construction, note that $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}\leq d-n<\infty.$ Next, note that $t=d$ is the only time point such that |$|\{x_{1:d}\}|=d$. Finally, note that when $s=t=d$, we have that $\mathcal{G}(x_{1:s})=\hat{x}\notin\operatorname{supp}(h)\setminus\{x_{1:s}\}$ by definition. If however $d<n$, then as soon as the common intersection $x_{1:n}$ is enumerated, the generator is forced to produce an output that necessarily lies outside of the support of some hypothesis in $\mathcal{H}.$ This completes the proof. ∎
3.2 Uniform Noise-dependent Generatability
Theorem 3.1 shows that obtaining guarantees that are uniform over the noise-level is generally hopeless. In this section, we study the easier setting of uniform noise-dependent generation, where the number of examples needed for perfect generation can depend on the noise-level. At a high level, the results in this section extend the techniques from Li et al. (2024) to account for noisy example streams. As such, we first define a noisy analog of the Closure dimension (see Appendix B for a definition).
Definition 3.2 ($n$-Noisy Closure dimension).
The Noisy Closure dimension of $\mathcal{H}$ at noise-level $n\in\mathbb{N}$, denoted $\operatorname{NC}_{n}(\mathcal{H})$, is the largest natural number $d$ for which there exist distinct $x_{1},\dots,x_{d}\in\mathcal{X}$ such that $\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},n}\neq\bot$ and $|\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},n}|<\infty.$ If this is true for arbitrarily large $d\in\mathbb{N}$, then we say that $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ On the other hand, if this is not true for $d=1$, we say that $\operatorname{NC}_{n}(\mathcal{H})=0.$
Unlike the Closure dimension, the Noisy Closure dimension is scale-sensitive, as it is defined with respect to every noise-level $n\in\mathbb{N}$. Scale-sensitive dimensions are not new to learning theory, but have also been defined and used to characterize other properties of hypothesis classes like PAC and online learnability for regression problems (Rakhlin et al., 2015; Bartlett et al., 1994). Our main theorem in this section uses the Noisy Closure dimension to provide a complete characterization of which classes are uniform noise-dependent generatable.
Theorem 3.3 (Characterization of Uniform Noise-dependent Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly noise-dependent generatable if and only if $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for all $n\in\mathbb{N}.$
Our proof of Theorem 3.3 is constructive. To prove necessity, we consider the case where there exists a $n\in\mathbbm{N}$ such that $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ Given any generator $\mathcal{G}$ and any number of distinct elements $d\in\mathbb{N}$, we explicitly pick an $h\in\mathcal{H}$ and a valid noisy stream $x_{1},x_{2},\dots$ such that $\mathcal{G}$ makes a mistake even after observing $d$ distinct examples. In fact, a simple modification of our proof shows that if $\operatorname{NC}_{n}(\mathcal{H})=d$, then any generator $\mathcal{G}$ must observe at least $d$ distinct examples before perfectly generating positive examples when the noise-level is $n$. To prove sufficiency, we explicitly construct a generator $\mathcal{G}$, which only needs to observe $\operatorname{NC}_{n}(\mathcal{H})+1$ distinct examples before being able to perfectly generate when the noise-level is $n\in\mathbb{N}$. Crucially, our generator $\mathcal{G}$ does not need to know the noise-level picked by the adversary. Together, our necessity and sufficiency directions show that not only does the finiteness of $\operatorname{NC}_{n}(\mathcal{H})$ at every $n\in\mathbbm{N}$ provide a qualitative characterization of uniform noise-dependent generatability, but it also provides a quantitative one – the “sample complexity” at noise-level $n\in\mathbbm{N}$ is $\Theta(\operatorname{NC}_{n}(\mathcal{H}))$ – analogous to the mistake bound in online learning, which quantifies learnability. We defer the full proof to Appendix E.
Our characterization of uniform noise-dependent generatability in terms of the Noisy Closure dimension allows us to show that all finite classes are uniformly noise-dependent generatable. This contrasts uniform noise-independent generatable, where even simple classes of size two may not be uniformly noise-independent generatable.
Corollary 3.4 (All Finite Classes are Uniformly Noise-dependent Generatable).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If $\mathcal{H}$ is finite, then $\mathcal{H}$ is uniformly noise-dependent generatable.
@@ 3.2 Uniform Noise-dependent Generatability
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be a finite hypothesis class with $q:=|\mathcal{H}|$ satisfying the $\operatorname{UUS}$ property. To show that $\mathcal{H}$ is uniformly noise-dependent generatable, it suffices to show that $\operatorname{NC}_{n}(\mathcal{H})<\infty$ for all $n\in\mathbb{N}.$ For any subset $V\subseteq\mathcal{H},$ define $\langle\emptyset\rangle_{V}:=\bigcap_{h\in V}\operatorname{supp}(h)$ and =$\mathcal{F}=\{V\subseteq\mathcal{H}:|\langle\emptyset\rangle_{V}|<\infty\}$ to be the set of examples common to all hypotheses in $V$ and the subsets of $\mathcal{H}$ whose intersection of supports has finite cardinality, respectively. Then, let $d:=\max_{V\in\mathcal{F}}|\langle\emptyset\rangle_{V}|$ be the maximum size of a finite set of examples common to a subset of $\mathcal{H}$. Note that $d$ is finite because $\mathcal{H}$ is finite. Fix any noise-level $n\in\mathbb{N}$ and for the sake of contradiction, suppose $\operatorname{NC}_{n}(\mathcal{H})=\infty.$ This means that for every $s\in\mathbb{N}$, there exists $t\geq s$ and a sequence of distinct examples ${x_{1},\dots,x_{t}}$ such that $|\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}|<\infty$. Pick $s=nq+d+1$ and consider the stream $x_{1},\dots,x_{t}$ such that $|\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}|<\infty$ for $t\geq s.$ Recall that $\mathcal{H}(x_{1:t};n)=\{h\in\mathcal{H}:|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h)|\geq t-n\}$. That is, $\mathcal{H}(x_{1:t};n)$ contains all hypotheses in $\mathcal{H}$ that are inconsistent with $x_{1:t}$ on at most $n$ examples. Since $|\mathcal{H}(x_{1:t};n)|\leq|\mathcal{H}|=q,$ there are at least $t-nq\geq(nq+d+1)-nq=d+1$ distinct examples in $x_{1:t}$ contained in the support of all hypotheses in $\mathcal{H}(x_{1:t};n)$. In other words, $|\langle x_{1},\dots,x_{t}\rangle_{\mathcal{H},n}|=|\langle\emptyset\rangle_{\mathcal{H}(x_{1:t};n)}|\geq d+1>d.$ This is a contradiction as $d=\max_{V\in\mathcal{F}}|\langle\emptyset\rangle_{V}|$ by definition and $\mathcal{H}(x_{1:t};n)\in\mathcal{F}.$ Thus, $\operatorname{NC}_{n}(\mathcal{H})<nq+d+1<\infty$ for all $n\in\mathbb{N},$ completing the proof. ∎
@@
We end this section by establishing that uniform noise-dependent generatability is strictly harder than noiseless uniform generatability. This is similar to online learnability for classification problems, where typically the noise-free (realizable) and noisy (agnostic) characterizations of learnability are not equivalent for deterministic learning algorithms (Littlestone, 1987; Ben-David et al., 2009).
Lemma 3.5 (Uniform Generatability $\neq$ Uniform Noise-dependent Generatability).
@@
Let =$\mathbb{P}=\{p_{1},p_{2},\dots\}$ be the set of primes, $S_{q}=\{-q^{n}:n\in\mathbb{N}\}$, $\mathbb{S}=\bigcup_{q\in\mathbb{P}}S_{q},$ and $\mathcal{X}=\mathbb{P}\cup\mathbb{S}$. Let the support of each hypothesis $h_{p}$ be $\operatorname{supp}(h_{p})=\mathbb{P}\setminus\{p\}\cup\mathbb{S}\setminus S_{p}$ and :=$\mathcal{H}:=\{h_{p}:p\in\mathbb{P}\}.$ Observe that $\mathcal{H}$ satisfies the $\operatorname{UUS}$ property and $\mathcal{X}$ is countable.
@@
We now prove that the Closure dimension (see Appendix B) $\operatorname{C}(\mathcal{H})=0.$ Suppose the first element in the sequence is $x_{1}=-p^{n}$ for some $p\in\mathbb{P}$ and $n\in\mathbb{N}.$ Then $|\langle x_{1}\rangle_{\mathcal{H}}|=|\{-p^{m}:m\in\mathbb{N}\ \}|=\infty.$ If the first element in the sequence is $x_{1}=p\in\mathbb{P},$ then $|\langle x_{1}\rangle_{\mathcal{H}}|=|\{-p^{n}:n\in\mathbb{N}\ \}|=\infty.$ This is because observing any example immediately rules out the one hypothesis that predicts 00 for that example. Note that these are the only two possible cases as it must be the case that $x_{1}\in\operatorname{supp}(h)$ for some $h\in\mathcal{H}.$
@@
Now, we show that $\operatorname{NC}_{1}(\mathcal{H})=\infty.$ Fix some $d\in\mathbb{N}.$ Consider the stream $p_{1},p_{2},\dots,p_{d}$ of the first $d$ prime numbers. Then, $|\langle p_{1},\dots,p_{d}\rangle_{\mathcal{H},1}|=|\emptyset|=0,$ due to the fact that $\mathcal{H}(p_{1},\dots,p_{d};1)=\mathcal{H}$ and the intersection of the supports of all $h\in\mathcal{H}$ is empty. So, $\operatorname{NC}_{1}(\mathcal{H})\geq d.$ Since $d\in\mathbb{N}$ was picked arbitrarily, this is true for all $d\in\mathbb{N}$ and therefore, $\mathcal{H}$ is not uniformly noise-dependent generatable. ∎
@@
Lemma 3.5 shows a strong separation between noise-free and uniform noise-dependent generation – there is a class that is uniformly generatable, but not uniformly noise-dependent generatable even when the adversary is allowed to perturb one example.
3.3 Non-uniform Noise-dependent Generatability